1500 lbs of brazilian coffee which sells at $3.80 per pound are mixed with 800 pounds of Jamacian Blue Mountain coffee which sells for $14.00 per pound. How should we price the blend if we mix the two amounts of coffee?
Answer by Maths68(1474)

1500 lbs of brazilian coffee which sells at $3.80 per pound are mixed with 800 pounds of Jamacian Blue Mountain coffee which sells for $14.00 per pound. How should we price the blend if we mix the two amounts of coffee?
Let
Brazillian Coffee = b
Jamacian Blue Mountain Coffee = j
Mixture = m
b price = $3.8
J Price = $14
m price = $x
b = 1500 lb
j = 800 lb
m = 1500+800 = 2300 lb
(b*price)+(j*price)=(m*price)
(3.8*1500)+(14*800)=2300*x
5700+11200=2300x
16900=2300x
16900/2300=x
7.38=x
x=7.38
Mixture will be sold at $7.38 per pound.